OpenStudy (anonymous):

find the average rate of change of f(x) = 3x^2 + 1 (a) from 2 to 4 (b) from -2 to 0 and (c) from 2 to 5

OpenStudy (anonymous):

Average rate of change is: \[ \frac{f(x_2)-f(x_1)}{x_2-x_1} \]I could derive it for you if you're curious.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Well we know that the rate of change is \(f'(x)\). So we want to find the average of that function. The average of a function is give by \[ g_{average} = \frac{1}{x_2-x_1}\int_{x_1}^{x^2}g(x)dx \]

OpenStudy (anonymous):

So we let \(g = f'\) and get: \[ f'_{average} = \frac{1}{x_2-x_1}\int_{x_1}^{x_2}f'(x)dx = \frac{f(x_2)-f(x_1)}{x_2-x_1} \]I just used the fundamental theorem of Calculus on that second step.

OpenStudy (anonymous):

Ok so how would I plug in the equation with the rate of change equation?

OpenStudy (anonymous):

I always mess that part up for some reason....

OpenStudy (anonymous):

(a) from 2 to 4 This means \(x_1 = 2,x_2=4\). So they want \[ \frac{f(4)-f(2)}{4-2} = \frac{(3(4)^2 + 1)-(3(2)^2 + 1)}{4-2} \]

OpenStudy (anonymous):

omg that is a lot

OpenStudy (anonymous):

So your replacing x for 4 and 2?

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

So how would I get the rate of change?